OpenStudy (anonymous):

Find the vertex of the graph of the function. f(x) = 4x2 + 24x + 32 (5, 4) (-3, -4) (-4, -3) (4, 5)

OpenStudy (primeralph):

Usually at -b/2a

OpenStudy (harsimran_hs4):

1. notice it`s an upward parabola 4(x^2 + 6x + 8) 4(x^2 + 2*3*x + 9) - 4 4(x+3)^2 -4 so when is value min : when x+3 is zero so x = -3 now find the value of this expression by putting x = -3 and get y coordinate

OpenStudy (anonymous):

Umm...I'm confused.

OpenStudy (harsimran_hs4):

which step ?

OpenStudy (anonymous):

I know you made the 3 negative but how can you get a y-coordinate?

OpenStudy (harsimran_hs4):

f(-3) = 4(-3)^2 +24(-3) + 32 = 36 - 72 + 32 = -4

OpenStudy (anonymous):

Duh. Oh okay. I get it! Thanks!!!!

OpenStudy (anonymous):

I forgot all about plugging it in the f(x). Thanks again!

OpenStudy (harsimran_hs4):

Welcome